# Tunability of the magnetic properties in Ni intercalated transition metal dichalcogenide NbSe2

### v1

@@ Acknowledgements.
The authors acknowledge A. Wadge, L. Plucinski, A. Wiśniewski, L. K. Tenzin, J. Sławińska and R. Sattigeri for useful discussions. This research was supported by the ”MagTop” project (FENG.02.01-IP.05-0028/23) carried out within the ”International Research Agendas” programme of the Foundation for Polish Science, co-financed by the European Union under the European Funds for Smart Economy 2021-2027 (FENG). C.A. acknowledges support from PNRR MUR project PE0000023-NQSTI. We further acknowledge access to the computing facilities of the Interdisciplinary Center of Modeling at the University of Warsaw, Grant g91-1418, g91-1419, g96-1808 and g96-1809 for the availability of high-performance computing resources and support. We acknowledge the access to the computing facilities of the Poznan Supercomputing and Networking Center, Grants No. pl0267-01, pl0365-01 and pl0471-01.Appendix A Computational details
We conducted calculations using the Vienna Simulation Package (VASP)Kresse and Hafner (1993); Kresse and Furthmüller (1996) based on density functional theory (DFT) by the projector-augmented wave (PAW) methodKresse and Joubert (1999). The exchange-correlation functional was described using the generalized gradient approximation (GGA) proposed by Perdew, Burke, and Ernzerhof (PBE)Perdew et al. (1996). To investigate the relationship between the ground state and the on-site Coulomb interaction U in detail, the DFT U method was involvedLiechtenstein et al. (1995), U was varied from 0 to 3 eV. Hund’s coupling JH was set to 0.15U, a value deemed suitable for 3 transition metal compounds. The k-point grid was set to for the primitive cell and for the supercell, with a total energy convergence criterion of eV and a cutoff energy of 450 eV. To calculate the magnetocrystalline anisotropy, we use a large k-grid for the unit cell. NiNb4Se8 crystallizes in the space group P63/ (no. 194). We need to double the lattice parameters in the plane using a 2 2 1 supercell; the lattice parameters of the supercell determined experimentallyWadge et al. (2025) become ==2=6.8889 Å and =12.54527 Å, this supercell is used to study the FM and AM phase. We need to further double the cell and construct the supercell 4 2 1 to study the AFM1 and AFM2 magnetic configurations, this last supercell holds the space group of P21. Lattice parameters are kept fixed while atomic positions are optimized for each magnetic structure.

### v2

@@ Acknowledgements.
The authors acknowledge A. Wadge, L. Plucinski, A. Wiśniewski, L. K. Tenzin, J. Sławińska, M. Cuoco, F. Mazzola, D. Romanin and M. Calandra for useful discussions. C. A. thanks R. Sattigeri for checking the sign of J|| for Ni0.25NbSe2 within Quantum Espresso. This research was supported by the ”MagTop” project (FENG.02.01-IP.05-0028/23) carried out within the ”International Research Agendas” programme of the Foundation for Polish Science, co-financed by the European Union under the European Funds for Smart Economy 2021-2027 (FENG). C.A. acknowledges support from PNRR MUR project PE0000023-NQSTI. We further acknowledge access to the computing facilities of the Interdisciplinary Center of Modeling at the University of Warsaw, Grant g91-1418, g91-1419, g96-1808 and g96-1809 for the availability of high-performance computing resources and support. We acknowledge the access to the computing facilities of the Poznan Supercomputing and Networking Center, Grants No. pl0267-01, pl0365-01 and pl0471-01.Appendix A Computational details
We conducted calculations using the Vienna Simulation Package (VASP)Kresse and Hafner (1993); Kresse and Furthmüller (1996) based on density functional theory (DFT) by the projector-augmented wave (PAW) methodKresse and Joubert (1999). The exchange-correlation functional was described using the generalized gradient approximation (GGA) proposed by Perdew, Burke, and Ernzerhof (PBE)Perdew et al. (1996). To investigate the relationship between the ground state and the on-site Coulomb interaction U in detail, the DFT U method was involvedLiechtenstein et al. (1995), U was varied from 0 to 3 eV. Hund’s coupling JH was set to 0.15U, a value deemed suitable for 3 transition metal compounds. We also performed tests using Local Density Approximation (LDA) or PAW with the 3p of Ni as semi-core to compare the AM and FM phase. In LDA, we found that the system is non-magnetic for U = 6 eV and below. Once the system becomes magnetic, we still obtain that the FM phase is the ground state. Including the 3p of Ni still does not alter the energy relationship between the FM and AM magnetic phases. The k-point grid was set to for the primitive cell of Ni0.25NbSe2 and Ni0.33NbSe2, for NiNbSe2, and for the supercell of Ni0.25NbSe2, with a total energy convergence criterion of eV and a cutoff energy of 450 eV. To calculate the magnetocrystalline anisotropy, we use a large k-grid for the unit cell. NiNb4Se8 crystallizes in the space group P63/ (no. 194). We need to double the lattice parameters in the -plane using a 2 2 1 supercell; the lattice parameters of the supercell determined experimentallyWadge et al. (2025b) become ==2=6.9100Å and =12.3238 Å, this supercell is used to study the FM and AM phase. We need to double the cell further and construct the supercell 4 2 1 to study the AFM1 and AFM2 magnetic configurations; this last supercell holds the space group of P21. Lattice parameters are kept fixed while atomic positions are optimized for each magnetic structure. The orbital magnetic moments are local moments on the Ni atoms calculated from the output of the VASP code. For weak ferromagnetism part, details regarding the relativistic spin density of state are reported in the literatureAutieri et al. (2025).